1. Solve for future value for each of the following.

Present Value $

Years

Interest Rate %

Future Value $

460

12

2.5

12,750

3

1.25

143,000

7.5

2.8

189,002

26

4.5
Answer #1

We can determine the fuuture value using the concept of time value of money. As per the concept of time value of money, the value of money available today is more compared to the same sum in some future time.
The formula used is:
Future value=Present value*(1+Interest rate)^(Number of years)

Future value=460*(1+2.5%)^(12)
=460*1.344888824
=$618.64886

Future value=12750*(1+1.25%)^(3)
=12750*1.037970703
=$13234.1265


Future value=143000*(1+2.8%)^(7.5)
=143000*1.230122494
=$175907.5166


Future value=189002*(1+4.5%)^(26)
=189002*3.140679007
=$593594.6137